OpenStudy (anonymous):

Find tan^-1 infinity/4.

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

is it \[\lim_{x\to \infty}\frac{\tan^{-1}(x)}{4}\]?

OpenStudy (anonymous):

No, tangent inverse of infinity/4.

OpenStudy (anonymous):

there is no such number as \(\frac{\infty}{4}\)

OpenStudy (anonymous):

as \(x\to \infty\) you have \(\tan^{-1}(x)\to \frac{\pi}{2}\) maybe that helps, not sure

OpenStudy (anonymous):

So the answer is pi/2 since that's the domain of tangent inverse?

OpenStudy (primeralph):

Probably.